# DiSCo Meets LLMs: A Unified Approach for Sparse Retrieval and Contextual Distillation in Conversational Search

### v1

@@ 4.1. Datasets and Metrics
Datasets. We evaluate our method’s effectiveness on various conversational passage retrieval datasets, namely:
@@
TREC CAsT 2020, 2022 (Dalton et al., 2020a; Owoicho et al., 2022) and TREC iKAT 2023 (Aliannejadi et al., 2024) are smaller conversational datasets, but carefully hand-crafted to include various conversational complexities. Moreover, they are coupled with rich query-passage relevance labels done by NIST assessors.
More details about the datasets can be found in Table 1.
@@
| TREC iKAT 23 | Test | 25 | 133 | 116M |
Metrics. We report the results in terms of main IR metrics used in various works (Yu et al., 2021; Mao et al., 2023c), as well as official TREC CAsT and iKAT reports (Dalton et al., 2020b; Aliannejadi et al., 2024). Our metrics are Mean Reciprocal Rank (MRR), Normalized Discounted Cumulative Gain (nDCG) (Järvelin and Kekäläinen, 2000) at 3. As we focus on first-stage retrieval, we also include Recall at different ranks (Xiao et al., 2023). We determine the statistically significant differences by doing a two-sided paired t-test with Bonferroni correction at 95% confidence ()
@@
where and are the probabilities of activation of the token in the vocabulary, resp. in query and document representations, over which we average on the dataset distribution333Note that this metric depends on the distribution of queries, thus the same model may have different FLOPs depending on the query input distribution..
We also report the number of LLM inference calls when used for query rewriting, as it is a strong indicator of inference latency. As an example, it takes an average of 4.4 seconds to generate 64 tokens on Llama3.1, on A100 GPU, while the typical dual-encoder retrieval latency is in the range of 100 milliseconds on CPU (Lassance and Clinchant, 2022a; Lassance et al., 2023, 2024a).
@@ 4.2. Baselines
Query rewriting methods. SPLADE-[T5/Llama/Mistral/Human]QR does retrieval using the SPLADE ad-hoc retrieval model trained on MS MARCO (Nguyen et al., 2016). As input, we pass the rewritten query using T5 (Raffel et al., 2019), Llama 3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), or gold human rewrites444We used the latest Llama meta-llama/Meta-Llama-3.1-8B-Instruct, Mistral mistralai/Mistral-7B-Instruct-v0.2 and T5 castorini/t5-base-canard. SPLADE no rewrite is the same ad-hoc retrieval model without rewrite, on the original conversation. IterCQR (Jang et al., 2024), CHIQ-Fusion (Mo et al., 2024a), and LLM4CS (Mao et al., 2023b) are state-of-the-art query rewriting baselines, however, they require multiple LLM calls at inference, which puts them under a high disadvantage.
Supervised fine-tuned methods. convSPLADE (Mao et al., 2023c) and convANCE (Mao et al., 2023c) are two methods fine-tuned using the InfoNCE loss on the conversational contrastive labels. Their input is the whole conversational context. We reproduce convSPLADE for out-of-domain retrieval using the same hyperparameters as the original paper.
Distillation-based method. LeCoRe (Mao et al., 2023c), QRACDR (Mo et al., 2024b) and ConvDR (Yu et al., 2021)555We did not manage to reproduce the results reported in the original papers, as our reproduced results were much lower. Therefore, we report the numbers in the original papers. Therefore, we are not able to conduct any significance tests with these models. all learn to distill the gold human rewrite representations and usually combine the distillation loss with an InfoNCE. Like our DiSCo, these models do not rely on LLM calls either, making them comparable. ConvDR and QRACDR are both dense approaches, using latent representations, while LeCoRe uses learned sparse representations. All of them distill query representations, while DiSCo the similarity scores, as a relaxation of their distillation method.
4.3. Implementations details
We use the SPLADE++ (Formal et al., 2022)666naver/splade-cocondenser-ensembledistil (Formal et al., 2022) checkpoints from Huggingface (Wolf et al., 2020) for all our SPLADE models. To further finetune on the CS task, we finetune for 5 epochs, with a learning rate of 2E-5, a max sequence length for queries of 64 tokens, and 100 tokens for answers, with a total limit of 256 tokens for the full conversational context. We use a batch size of 10, with 16 negatives per query, and in-batch negatives (Lin et al., 2021b). Default experiments with SPLADE use FLOPS and L1 regularization, as in the original code, with values , . We use mixed precision and fp16 to maximize the use of the memory. For sparse retrieval, we use inverted indexes based on the numba library and pyserini (Lin et al., 2021a) together with Pytorch (Paszke et al., 2019). The fusion method uses the ranx library (Bassani, 2022). All teachers use In-Context learning to generate the rewrites. We list the prompts to generate the rewrites in Appendix C.
Given the high efficiency of our fine-tuning, we were able to run the experiments on a single A100 GPU with 40GB of GPU memory.
@@ 4.3. Implementations details
| SPLADE T5QR | T5 | 1 | 0.667 | 0.501 | 0.321 | 0.314 | 0.840 | 0.617 | 0.382 | 0.366 | ||
| SPLADE LlamaQR () | Llama 3.1 | 1 | 0.761 | 0.572 | 0.365 | 0.352 | 0.826 | 0.613 | 0.377 | 0.360 | ||
@@ 5. Results
5.1. Performance Comparison
In this first subsection, we aim to answer RQ1 and RQ2 by comparing the performance of our proposed DiSCo with state-of-the-art query rewriting, and representation distillation methods. We first compare the performance of DiSCo in the in-domain and out-of-domain retrieval setups in Tables 2 and 3, respectively. In the table, we report the performance of diverse baselines in terms of various metrics, as well as the number of LLM calls every model requires at inference. This is an important factor while comparing the performance of conversational retrieval models for various reasons, namely, (i) LLM call leads to considerable inference latency, delaying inferencing from hundreds of milliseconds to seconds; and (ii) LLM-based methods take advantage of the vast parameter size and knowledge learned by the training of the LLMs, making their comparison to other smaller models unfair.
In-domain retrieval. To answer RQ1, we see in Table 2 that in the in-domain retrieval setting on TopiOCQA and QReCC, we do not observe a big gap between rewriting and distillation-based method, even though the rewriting-based methods make use of the LLM knowledge in the rewriting phase. Generally, distillation-based methods (i.e., ConvDR, QRACDR, LeCoRe) outperform the rewriting-based methods while being more efficient too as they have zero LLM calls. This indicates that the distillation phase already knows the collection and topic of the queries effectively.
Out-of-domain retrieval. To further study the effectiveness of our proposed distillation approach, and study its generalizability, in Table 3 we report the results in the out-of-domain retrieval setting on TREC CAsT 20,22 and iKAT 23. Addressing RQ2, we see that, unlike the in-domain setting, there is a considerable gap between the rewriting-based and distillation-based methods when it comes to out-of-domain retrieval. On average, rewriting-based methods perform 19% better than distillation-based methods because they mainly rely on the massive parameter size and knowledge of LLMs, leading to high generalizability. However, as mentioned earlier making LLM calls on inference puts the models at a high disadvantage because of the high latency.
Focusing on the distillation-based approaches, our DiSCo outperforms other methods by a large margin, showing the remarkable generalizability of our proposed relaxed distillation. In particular, we see that DiSCo Human outperforms QRACDR by 3.8 points in terms of nDCG3, and LeCoRe by 5.5 points in terms of Recall@100, both on CAsT 2020. Note that those two models are two state-of-the-art models in the same zero-shot settings, trained on QReCC and evaluated on CAsT 2020. It is also noteworthy that even though our distillation-based approach is not able to outperform some of the rewriting-based (LLM-based) methods, our DiSCo-Fusion, which is a fusion method based on LLM rewriting and distillation manages to outperform all the rewriting-based methods by a large margin in terms of Recall@100, reaching SotA conversational passage retrieval performance. When comparing closely with the SotA LLM4CS model, we note that while LLM4CS takes advantage of 5 GPT-3.5 calls, our model outperforms it in terms of recall by only one Mistral LLM call. Comparing the precision-oriented metrics, we see that LLM4CS outperforms DiSCo-Fusion, on the cost of 12 recall points. Considering that we focus on the retrieval task, we consider recall to be preferable, as reranking can be added post-retrieval on a smaller set of documents (Xiao et al., 2023; Déjean et al., 2024).
@@ 5.1. Performance Comparison
| R@100 | MRR | nDCG3 | R@100 | MRR | nDCG3 | R@100 | MRR | nDCG3 | ||||||
@@
| SPLADE LlamaQR | Llama 3.1 | 1 | 0.550 | 0.515 | 0.376 | 0.312 | 0.453 | 0.300 | 0.198 | 0.281 | 0.177 | |||
@@
Distillation-rewriting fusion. Furthermore, note in Table 3 that DiSCo-Fusion is the fusion of our DiSCo and SPLADE MistralQR, as the fusion of the student with the teacher. Even though SPLADE MistralQR exhibits high performance (e.g., 0.57 and 0.33 in terms of Recall@100 resp. on CAsT 2020, 2022), fusing it with the distilled DiSCo leads to further significant improvements (4 points increase on both datasets). This indicates that relaxing the distillation process to similarity score enables the model to go beyond the knowledge of the teacher in the representation space. This gain is even stronger on the precision of CAsT 2022, where DiSCo-Fusion outperforms SPLADE MistralQR by 8 points in terms of nDCG3.
@@
Comparison with the teacher models. Now looking at the performance of the students and teachers models on Table 2, we see that the student models outperform the teachers significantly for in-domain retrieval. In particular, DiSCo Human outperforms SPLADE HumanQR by 3.5 MRR points on QReCC; and similarly with DiSCo Mistral. This demonstrates that our distillation objective allows the student model to learn representations that surpass the original teacher’s representations. This is possible thanks to the relaxation of the distillation, as the student can adapt representations. This was not possible with the previous distillation objective on the representation, as student models were trained to copy the representations of the teacher model. Out-of-domain however, we see in Table 3 that the student models have more difficulties outperforming the performance of the teacher models. This result on out-of-domain retrieval is however not fair considering that teacher model generalization relies on billions of parameters, while students on a few hundred. Recall also that DiSCo-Fusion as the fusion of the teacher and the student was further improving performance out-of-domain.
@@ 5.2. Multiple Teachers Distillation
While the previous section focused on the distillation of single teachers, we try here to answer RQ3 on the use of multiple teachers. We thus compare our DiSCo model trained with multiple teachers.
In-domain and out-of-domain retrieval. Trying to address RQ3, both Tables 2 and 3 include our DiSCo multi-teach model trained with multiple teachers. Considering the performances of DiSCo multi-teach for in-domain retrieval, we observe significant gains compared to DiSCo Mistral on QReCC. In particular, we observe a 1-point increase in terms of R@10, MRR, and nDCG3 on QReCC. This result remains consistent for out-of-domain retrieval, where within Table 3 DiSCo multi-teach outperforms its single version model, by at least 2 points on all metrics, on CAsT 2020 and 2022. On iKAT 2023 however, we observe mixed results. Two interpretations are possible for these general gains: first, the multiple teachers and their subsequent rewrites produce word synonyms, as a form of query expansion, improving the performance, second, as one LLM may fail to resolve the context modeling task, another can balance it, offering better robustness.
Progressive Improvement. Further addressing RQ3, Figure 4 gives insights on the multiple teachers distillation. While the left part of the Figure gives the performances of the teachers on QReCC, the right part shows the gains when incrementally adding teachers to the distillation. We can also notice the important gains from the distillation of LlamaQR. In particular, DiSCo Llama improves by 8 points compared to SPLADE LlamaQR in terms of Recall@100, while in comparison DiSCo Mistral by only 4 points compared to his teacher. This shows that the relaxation is robust, and can learn from even lower quality rewrite.
This demonstrates that we can benefit from multiple LLM teachers during distillation, with benefits on the generalization capacities of the student. This answers RQ3, on the knowledge distillation from multiple teachers.

### v2

@@ 4.1. Datasets and Metrics
Datasets. We evaluate our method’s effectiveness on various conversational passage retrieval datasets, detailed in Table 1, namely:
@@
TREC CAsT 2020, 2022 (Dalton et al., 2020a; Owoicho et al., 2022) and TREC iKAT 2023 (Aliannejadi et al., 2024) are smaller conversational datasets, but carefully hand-crafted to include various conversational complexities. They are coupled with high-quality relevance judgements done by NIST assessors.
@@
| TREC iKAT 23 | Test | 25 | 176 | 116M |
Effectiveness Metrics. We report the results in terms of the main IR metrics (Yu et al., 2021; Mao et al., 2023c), as well as official TREC CAsT and iKAT ones (Dalton et al., 2020b; Aliannejadi et al., 2024). Metrics are Mean Reciprocal Rank (MRR), Normalized Discounted Cumulative Gain (nDCG) (Järvelin and Kekäläinen, 2000) at 3. As we focus on first-stage retrieval, we also include Recall at different ranks (Xiao et al., 2023). We determine the statistically significant differences by doing a two-sided paired t-test with Bonferroni correction at 95% confidence ().
Efficiency Metrics. In terms of inference efficiency, we report the number of LLM inference calls when used for query rewriting, as it is a strong indicator of inference latency. As an example, it takes an average of 4.4 seconds to generate 64 tokens on Llama 3.1, on A100 GPU, while the typical dual-encoder retrieval latency is in the range of 100 milliseconds on CPU (Lassance and Clinchant, 2022a; Lassance et al., 2023, 2024a).
We also provide Rewrite and Retrieval efficiency, of several models. This is performed per query, averaged over the entire test set. Rewrite is the latency for rewriting, while retrieval is both query encoding and search through the inverted index (in the case of SPLADE models). We used a 4th AMD EPYC CPU and an A100 GPU.
@@
where and are the probabilities of activation of the token in the vocabulary, resp. in query and document representations, over which we average on the dataset distribution.
@@ 4.2. Baselines
Query rewriting methods. SPLADE-[T5/Llama/Mistral/Human]QR does retrieval using the SPLADE ad-hoc retrieval model trained on MS MARCO (Nguyen et al., 2016). As input, we pass the rewritten query using T5 (Raffel et al., 2019), Llama 3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), or gold human rewrites222We used the latest Llama meta-llama/Meta-Llama-3.1-8B-Instruct, Mistral mistralai/Mistral-7B-Instruct-v0.2 and T5 castorini/t5-base-canard. SPLADE no rewrite is the same ad-hoc retrieval model without rewrite, on the original conversation. IterCQR (Jang et al., 2024), CHIQ-Fusion (Mo et al., 2024a), and LLM4CS (Mao et al., 2023b) are state-of-the-art query rewriting baselines; however, they require multiple LLM calls at inference, which puts them under a high disadvantage. For LLM4CS, we reproduced their best setting (RAR, Mean aggregation from , CoT, GPT-4).
Supervised fine-tuned methods. convSPLADE (Mao et al., 2023c) and convANCE (Mao et al., 2023c) are two methods fine-tuned using the InfoNCE loss on the conversational contrastive labels. Their input is the whole conversational context. We reproduce convSPLADE for out-of-domain retrieval using the same hyperparameters as in the original paper.
Distillation-based method. LeCoRe (Mao et al., 2023c), QRACDR (Mo et al., 2024c) and ConvDR (Yu et al., 2021)333We did not manage to reproduce the results reported in the original papers, as our reproduced results were much lower. Therefore, we report the numbers in the original papers. Therefore, we are not able to perform any significance tests with these models. all learn to distill the gold human rewrite representations and usually combine the distillation loss with an InfoNCE. Like our DiSCo, these models do not rely on LLM calls either, making them comparable. ConvDR and QRACDR are both dense approaches, using latent representations, while LeCoRe uses learned sparse representations. All of them distill query representations, while DiSCo the similarity scores, as a relaxation of their distillation method.
@@
| SPLADE T5QR () | T5 | 1 | 0.667 | 0.501 | 0.321 | 0.314 | 0.840 | 0.617 | 0.382 | 0.366 | ||
| SPLADE LlamaQR () | Llama 3 | 1 | 0.761 | 0.572 | 0.365 | 0.352 | 0.826 | 0.613 | 0.377 | 0.360 | ||
@@
4.3. Implementations details
We use the SPLADE++ (Formal et al., 2022)444naver/splade-cocondenser-ensembledistil (Formal et al., 2022) checkpoints from Huggingface (Wolf et al., 2020) for all our SPLADE models. To further finetune on the CS task, we fine-tune for 5 epochs, with a learning rate of , a max sequence length for queries of 64 tokens, and 100 tokens for answers, with a total limit of 256 tokens for the full conversational context. We also use a 256-token limit for passages. We use a batch size of 10, with 16 negatives per query, and in-batch negatives (Lin et al., 2021b). Default experiments with SPLADE use FLOPS and L1 regularization, as in the original code, with values , . We use mixed precision and fp16 to maximize memory use. For sparse retrieval, we use inverted indexes based on the numba library and pyserini (Lin et al., 2021a) together with Pytorch (Paszke et al., 2019). The fusion method uses the ranx library (Bassani, 2022). All teachers use In-Context learning to generate the rewrites. We provide the one-shot prompt used for rewriting with the LLM Teacher models in Table 2. We experimented in both zero-shot and one-shot, but decided to use one-shot to improve the quality of the teacher models. Given the high efficiency of our fine-tuning, we were able to run the experiments on a single A100 GPU with 40GB of GPU memory.
@@ 5. Results
| R@100 | MRR | nDCG@3 | R@100 | MRR | nDCG@3 | R@100 | MRR | nDCG@3 | ||||||
@@
| SPLADE LlamaQR | Llama 3 | 1 | 0.550 | 0.515 | 0.376 | 0.312 | 0.453 | 0.300 | 0.198 | 0.281 | 0.177 | |||
@@
5.1. Performance Comparison
In this first subsection, we aim to answer RQ1 and RQ2 by comparing the performance of our proposed DiSCo with state-of-the-art query rewriting, and distillation-based methods. We first compare the performance of DiSCo in the in-domain and out-of-domain retrieval setups in Tables 3 and 4, respectively. The tables report the performance of diverse baselines in terms of various metrics, as well as the number of LLM calls every model requires at inference. This is an important factor while comparing the performance of conversational retrieval models for various reasons, namely, (i) LLM call leads to considerable inference latency, delaying inferencing from hundreds of milliseconds to seconds; and (ii) LLM-based methods take advantage of the vast parameter size and knowledge learned by the training of the LLMs, making their comparison to other smaller models unfair.
In-domain retrieval. Trying to address RQ1, we report in Table 3 in-domain performance on TopiOCQA and QReCC. Looking at the results, we do not observe a big gap between the rewriting and distillation-based methods, even though the rewriting-based methods make use of the LLM knowledge in the rewriting phase. Distillation-based methods (i.e., ConvDR, QRACDR, LeCoRe) even outperform the rewriting-based methods while being more efficient too, as they have zero LLM calls.
Out-of-domain retrieval. To further study the effectiveness of our proposed distillation approach, and study its generalizability, we report in Table 4 the results in the out-of-domain retrieval setting on TREC CAsT 20,22 and iKAT 23. Addressing RQ2, we see that, unlike the in-domain setting, there is a considerable gap between the rewriting-based and distillation-based methods when it comes to out-of-domain retrieval. On average, rewriting-based methods perform 19% better than distillation-based methods because they mainly rely on the massive parameter size and knowledge of LLMs, leading to high generalizability. However, as mentioned earlier making LLM calls on inference puts the models at a high disadvantage because of the high latency.
Focusing on distillation-based approaches, our DiSCo outperforms other methods by a large margin, showing the remarkable generalizability of our proposed relaxed distillation. In particular, we see that DiSCo Human outperforms QRACDR by 3.8 points in terms of nDCG@3, and LeCoRe by 5.5 points in terms of Recall@100, both on CAsT 2020. Note that those two models are two state-of-the-art models in the same zero-shot settings, trained on QReCC and evaluated on CAsT 2020. It is also noteworthy that even though our distillation-based approach is not able to outperform some of the rewriting-based (LLM-based) methods, our DiSCo-Fusion, which is a fusion method based on LLM rewriting and distillation manages to outperform all rewriting-based methods by a large margin in terms of Recall@100, reaching SotA conversational passage retrieval performance. When comparing closely with the SotA LLM4CS model, we note that while LLM4CS takes advantage of 5 GPT-4 calls, our DiSCo multi-teach model outperforms it in terms of recall on all out-of-domain datasets. Comparing the precision-oriented metrics, we see that DiSCo multi-teach outperforms LLM4CS on all datasets, except TREC CAsT 2020. This dataset was the second edition of TREC CAsT and is considered to be simpler than both 2022 and 2023 versions, showing the performance of our model on complex information needs. Also, considering that we focus on the retrieval task, we consider recall to be preferable, as reranking can be added post-retrieval on a smaller set of documents (Xiao et al., 2023; Déjean et al., 2024).
@@ 5.1. Performance Comparison
Distillation-rewriting fusion. Furthermore, note in Table 4 that DiSCo-Fusion is the fusion of our DiSCo and SPLADE MistralQR, as the fusion of the student with the teacher. Although SPLADE MistralQR exhibits high performance (e.g., 0.57 and 0.33 in terms of Recall@100 resp. on CAsT 2020, 2022), fusing it with the distilled DiSCo leads to further significant improvements (4 points increase on both datasets). This indicates that relaxing the distillation process to the similarity score enables the model to go beyond the knowledge of the teacher in the representation space. This gain is even stronger on the precision of CAsT 2022, where DiSCo-Fusion outperforms SPLADE MistralQR by 8 points in terms of nDCG@3.
@@
Comparison with the teacher models. Now looking at the performance of the students and teachers models in Table 3, we see that the student models outperform the teachers significantly for in-domain retrieval. In particular, DiSCo Human outperforms SPLADE HumanQR by 3.5 MRR points on QReCC; and similarly with DiSCo Mistral. This demonstrates that our distillation objective allows the student model to learn representations that surpass the original teacher’s representations. This is possible thanks to the relaxation of the distillation, as the student can adapt and learn optimal representations based on gradient descent optimization. This was not possible with the previous distillation objective on the representation, as student models were trained to copy the representations of the teacher model. Out-of-domain however, we see in Table 4 that the student models have more difficulties outperforming the performance of the teacher models. This result on out-of-domain retrieval is, however, not fair considering that teacher model generalization relies on billions of parameters, while students on a few hundred. Recall also that DiSCo-Fusion as the fusion of the teacher and the student was further improving performance out-of-domain.
Reliance on weak teacher model. Finally, we explore the reliance on the teacher models, by training DiSCo with a weaker teacher model, such as T5. This is interesting in scenarios where we do not have access to strong teachers and also shows the robustness of the method when trained on more noisy labels. From Table 3, we see that DiSCo T5 performs almost on par with DiSCo Mistral, while SPLADE MistralQR outperforms SPLADE T5QR by an important margin, on both TopiOCQA and QReCC. Our method is thus robust even when trained on lower-quality rewrites from the teacher.
@@ 5.2. Multiple Teachers Distillation
While the previous section focused on the distillation of single teachers, here we answer RQ3 on the use of multiple teachers. Thus, we compare our DiSCo model trained with multiple teachers.
In-domain and out-of-domain retrieval. Trying to address RQ3, both Tables 3 and 4 include our DiSCo multi-teach model trained with multiple teachers. Considering the performances of DiSCo multi-teach for in-domain retrieval, we observe significant gains compared to DiSCo Mistral on QReCC. In particular, we observe a 1-point increase in terms of R@10, MRR, and nDCG@3 on QReCC. This result remains consistent for out-of-domain retrieval, where within Table 4 DiSCo multi-teach outperforms its single version model, by at least 2 points on all metrics, on CAsT 2020 and 2022. On iKAT 2023, however, we observe mixed results. Two interpretations are possible for these general gains: first, the multiple teachers and their subsequent rewrites produce word synonyms, as a form of query expansion, improving the performance, and second, as one LLM may fail to resolve the context modeling task, another can balance it, offering better robustness.
| Efficiency (ms) | Rewrite | Retrieval | Total | R@100 |
| SPLADE T5QR | 190 | 358 | 548 | 0.479 |
| SPLADE LlamaQR | 4245 | 367 | 4612 | 0.550 |
| SPLADE MistralQR | 2094 | 356 | 2450 | 0.572 |
| convSPLADE | 0 | 356 | 356 | 0.446 |
| DiSCo Mistral | 0 | 346 | 346 | 0.519 |
| DiSCo multi-teach | 0 | 357 | 357 | 0.531 |
| DiSCo Fusion | 2094 | 356 | 2450 | 0.611 |
Progressive Improvement. Further addressing RQ3, Figure 4 gives insights on the multiple teachers distillation. While the left part of the Figure gives the performances of the teachers on QReCC, the right part shows the gains when incrementally adding teachers to the distillation. We can also notice the important gains from the distillation of LlamaQR. In particular, DiSCo Llama improves by 8 points compared to SPLADE LlamaQR in terms of Recall@100, while in comparison DiSCo Mistral by only 4 points compared to his teacher. This shows that the relaxation is robust and can learn from even lower-quality rewrites.